OpenStudy (anonymous):

Hello! How would you solve Log_5(2x)=log_5(2^3)-log_5(12)?

jimthompson5910 (jim_thompson5910):

First you'll use log rule #2 described on this page http://www.purplemath.com/modules/logrules.htm

jimthompson5910 (jim_thompson5910):

That will turn \[\Large \log_{5}(2x) = \log_{5}(2^3) - \log_{5}(12)\] into \[\Large \log_{5}(2x) = \log_{5}\left(\frac{2^3}{12}\right)\]

jimthompson5910 (jim_thompson5910):

agreed so far?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

we have logs of the same base on either side of the equal sign so the stuff inside the logs must be equal \[\Large \log_{5}(2x) = \log_{5}\left(\frac{2^3}{12}\right)\] \[\Large 2x = \frac{2^3}{12}\] solve for x

OpenStudy (anonymous):

well thats the thing. My answer key provided by my teacher said that x was 1/3 but i keep getting 3

OpenStudy (anonymous):

oh wait! I got 1/3. Thank you so much!

jimthompson5910 (jim_thompson5910):

glad to be of help